OpenStudy (anonymous):

helpppp meeee :) 88x^3-16x^2+8x-11/8x

OpenStudy (nottim):

i think you could factor x out first

OpenStudy (anonymous):

yea but then what ---ugh

OpenStudy (nottim):

then its just a trinomial right? you could do complex trinomials, quadratics or simple trinomials.

OpenStudy (stacey):

What do the directions say to do with this?

OpenStudy (anonymous):

Is the problem\[\frac{88x^3-16x^2+8x-11}{8x}=11x^2-2x+1-\frac{11}{8x}\]

OpenStudy (anonymous):

i am so stupid...i just dont get it!! ugh

OpenStudy (anonymous):

animal love u!!!

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

math great...algebra not so great!! lol

OpenStudy (anonymous):

nothing is jumping out at me but stress

OpenStudy (anonymous):

Note that if you write the problem down correctly, the answer just about jumps out at you.

OpenStudy (anonymous):

Stay after it....

OpenStudy (anonymous):

so now I can cancel the 8x out and left with 11x^2-2x-10 ?

OpenStudy (anonymous):

Not exactly. The last term has no way to divide by 8x, so the denominator stays. Let me write it a little differently to give you a better view of it.

OpenStudy (anonymous):

\[\frac{88x^3-16x^2+8x-11}{8x}=\frac{88x^3}{8x}+\frac{-16x^2}{8x}+\frac{8x}{8x}+\frac{-11}{8x}=11x^2-2x+1-\frac{11}{8x}\]

OpenStudy (anonymous):

Does that help you see it?

OpenStudy (anonymous):

so I am basically factoring an 8x out?

OpenStudy (anonymous):

Exactly, except the last term.

OpenStudy (anonymous):

aaaahhhh yes

OpenStudy (anonymous):

Hope I was helpful.

OpenStudy (anonymous):

yes u were!!!! wanna help me with another!!???

OpenStudy (anonymous):

Sure, what else would I do? lol

OpenStudy (anonymous):

i wish i had a good grasp @ algebra.... ugh

OpenStudy (anonymous):

Not impossible. Just stay after it.

OpenStudy (anonymous):

I try... ok here goes.... 21x^3-10x^2+32x 48/7x+6

OpenStudy (anonymous):

Confirm the question. I think you left out an addition sign. Is it\[\frac{21x^3-10x^2+32x 48}{7x+6}\]

OpenStudy (anonymous):

+48

OpenStudy (anonymous):

see not good @ this lol

OpenStudy (anonymous):

Of course... I figured that out, then didn't fix it when I typed.... This one will need long division.

OpenStudy (anonymous):

I can do this problem, and even tell you how to do it, but I'm not very good at the draw function to show you the work. The structure of the problem exceeds my computer skill. Let me find a tutorial that will show you how. Wait a sec.....

OpenStudy (anonymous):

thanks so so much this bites lol

OpenStudy (anonymous):

cool..thanks

OpenStudy (anonymous):

Another that is pretty nice. http://tutorial.math.lamar.edu/Classes/Alg/DividingPolynomials.aspx

OpenStudy (anonymous):

wow thanks so much....it breaks down nicely

OpenStudy (anonymous):

This list is on my teaching website, it has some sites that might be helpful. http://complabs.nevada.edu/~ain/help%20sites.htm

OpenStudy (anonymous):

thank you so so much...u have been awesome!!!

OpenStudy (anonymous):

No sweat. Did I remind you to do math every day?

OpenStudy (anonymous):

yes I will =)... really have no choice...taking it in college

OpenStudy (anonymous):

I get it. Remember, if you're not having fun, you're not doing it right.

OpenStudy (anonymous):

uhmm ok lol....have good nite